### v1

@@ 1. Introduction
The theory of positive linear maps plays a central role in operator algebras and quantum information theory [Stø63, Cho75, Wor76, Pau02, Wat18]. Among the various refinements of positivity, the notion of $k$-positivity has emerged as a fundamental yet notoriously subtle concept. A linear map $\mathcal{L}:M_{d_{A}}(\mathbb{C})\to M_{d_{B}}(\mathbb{C})$ between matrix algebras is said to be $k$-positive if its ampliation map $\mathrm{id}_{k}\otimes\mathcal{L}:M_{k}(\mathbb{C})\otimes M_{d_{A}}(\mathbb{C})\to M_{k}(\mathbb{C})\otimes M_{d_{B}}(\mathbb{C})$ maps positive semidefinite operators to positive semidefinite operators. While complete positivity ($k=\min(d_{A},d_{B})$) admits several powerful characterizations, including the Choi–Kraus representation, the structure of $k$-positive maps for intermediate values of $k$ remains poorly understood. In fact, even deciding ($1$-)positivity is computationally hard [Gur03, GHP10], and several semidefinite programming relaxations for testing $k$-positivity have been proposed recently [CCF25]. Nevertheless, no general structural classification is known beyond very special cases.
@@
In the case of the $d\otimes d$ system, evidence for the existence of PPT states Schmidt number scales linearly in $d$ was first obtained in [SWZ11] using tools from asymptotic geometric analysis, although explicit constructions remained elusive. Early explicit examples exhibited PPT states whose Schmidt number grows logarithmically in $d$ [CYT17], while subsequent works improved this bound to linear scaling in $d$ [HLLMH18, PV19, Car20]. At present, the best known general upper bound is $\lfloor d/2\rfloor$, except for the dimensions $d=3$ [Cho82, Hor97] and $d=5$ [KG24], and explicit constructions achieving Schmidt number of order $d/2$ has been known only in limited settings [PV19, Car20]. The well-known dual formulation of this problem asks for the existence of $k$-positive indecomposable (i.e., not of the form $\mathcal{L}_{1}+\top\circ\mathcal{L}_{2}$ for some completely positive maps $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$) linear maps on $M_{d}(\mathbb{C})$ as large as possible. This dual viewpoint has also driven much of the development of the theory of $k$-positive maps and its Choi matrices [CK09, YLT16, HLLMH18, RBO20, MOM25, HK25a, HK25b, vEKD25].
@@
In this work, we contribute to both of the above questions on $k$-positivity and Schmidt numbers by exploring new classes of quantum objects possessing symmetry with respect to the symplectic group. More precisely, we study linear maps that are covariant under conjugation by unitary symplectic matrices, as well as bipartite quantum states that are invariant under $S\otimes S$ or $S\otimes\overline{S}$ actions, each parametrized by two real variables. The representation theory of the symplectic group gives rise to algebraic structures that are formally similar to those arising in symmetry classes associated with the orthogonal group [VW01, PY24]; however, the resulting entanglement properties turn out to be drastically different. As we show, this symmetry setting supports a remarkably rich entanglement landscape, including highly entangled PPT states and highly positive yet indecomposable linear maps.
@@
    in a $d\otimes d$ system, whenever it is well defined, is computed explicitly for all values of $(a,b)$. Furthermore, there exists a (triangular) set $S=S^{(d)}\subset\mathbb{R}^{2}$ with nonempty interior such that $\rho_{a,b}^{(d)}$PPT with the Schmidt number $d/2$’ if and only if $(a,b)\in S$, while ${\rm SN}\big(\rho_{a,b}^{(d)}\big)\leq d/2$ whenever $\rho_{a,b}^{(d)}$ is PPT.
@@
  • When $p=q=-\frac{1}{d-2}$, the linear map $\mathcal{L}_{p,q}^{(d)}$ corresponds to the widely studied Breuer-Hall map [Bre06, Hal06] $\mathcal{L}^{\rm BH}$, which has proven effective in detecting PPT entanglement. We further show, however, that $\mathcal{L}^{\rm BH}$ is not 2-positive, meaning it cannot serve as an entanglement witness for states with Schmidt number strictly larger than 2.
@@
    ${\rm SN}(\rho)-{\rm SN}(\rho^{\Gamma})=d/2-2.$
@@
In particular, we emphasize a new construction of Breuer–Hall–type maps, which are shown to be optimal $k$-positive indecomposable maps and thus remedy a limitation of the original Breuer–Hall map.
Theorem 1.2 (Theorem 3.7).
Let $d\geq 4$ be an even integer and let $V$ be a skew-symmetric unitary matrix. Then a linear map
@@
is an optimal $k$-positive indecomposable linear map for every $k=1,\ldots,d/2-1$. In other words, there exists no other $k$-positive witness that is able to detect strictly more PPT states with Schmidt number larger than $k$.
@@
We have two additional applications using our framework. In Section 5 we show that the PPT squared conjecture [Chr12, CMHW19] holds within symplectic group symmetry in two perspectives: (1) any composition of two PPT maps is entanglement-breaking, and (2) any composition of a positive map and a PPT map is decomposable. This result is in fact a simple consequence of Theorem 1.1. Nevertheless, this is still remarkable in the sense that the verification applies to a broad class of quantum channels whose Choi matrices may exhibit high-dimensionality of entanglement, thereby providing new and nontrivial evidence in support of the conjecture.
@@ 2.3. Symplectic group symmetry
$V=\Omega_{d}:=\begin{pmatrix}0&I_{d/2}\\ -I_{d/2}&0\end{pmatrix},\quad\langle x,y\rangle_{\Omega}:=\sum_{i=1}^{d/2}(x_{i}y_{i+d/2}-x_{i+d/2}y_{i}).$
@@ 3.1. Positivity, complete positivity, and decomposability
Now suppose $\mathcal{L}_{p,q}$ is decomposable, and write $\mathcal{L}_{p,q}=\mathcal{L}^{(1)}+\mathcal{L}^{(2)}$ where $\mathcal{L}^{(1)}$ is CP and $\mathcal{L}^{(2)}$ is co-CP. By applying the $(S,S)$-twirling operation $\mathcal{T}_{S,S}:=\mathcal{T}_{\iota_{Sp(d)},\iota_{Sp(d)}}$, we have $\mathcal{L}_{p,q}=\mathcal{T}_{S,S}\mathcal{L}^{(1)}+\mathcal{T}_{S,S}^{(2)}$ where $\mathcal{T}_{S,S}\mathcal{L}^{(1,2)}$ are $(S,S)$-covariant CP and is co-CP maps, resp, by Proposition 2.5. Therefore, the trace-preserving property of $\mathcal{L}_{p,q}$, combined with Proposition 3.1, implies that
@@ 3.2. $k$-Positivity
It is straightforward to verify that the point $P=(-\frac{2}{d^{2}-3d},-\frac{d-2}{d^{2}-3d})$ satisfies the inequalities above, and hence $P\in\mathbb{P}_{d/2-1}\setminus\mathbb{D}$. On the other hand, On the other hand, the boundary line segment of $\mathbb{P}_{d/2}$ connecting the vertices $(-\frac{2}{d^{2}-d-2},-\frac{d}{d^{2}-d-2})$ and $(-\frac{2}{d^{2}-2},0)$ lies entirely within the half-plane $x+y\geq-\frac{2+d}{d^{2}-d-2}$. This implies $\mathbb{P}_{d/2}\subset\mathbb{D}$. ∎
@@
We recover the result of [Bre06] in the case $k=1$, in a concise manner by exploiting symplectic group symmetry. We also remark that, without imposing the indecomposability condition, the same method allows us to characterize a class of optimal $k$-positive linear maps within the class ${\rm Cov}(S,S)$. For example, for $k=1,\ldots,d$, the $k$-reduction map
@@
is optimal in the sense that there exists no nonzero CP map $\Phi$ such that $\mathcal{L}^{\prime}=\mathcal{L}_{\frac{-1}{kd-1},0}-\Phi$ remains $k$-positive, as also established in [SBL01].
@@ 5. The PPT squared conjecture under symplectic group symmetry
    If $\mathcal{L}$ is a positive map that is either $(S,S)$- or $(S,\overline{S})$-covariant, and $\mathcal{L}_{2}$ is a PPT map that is either $(S,S)$- or $(S,\overline{S})$-covariant, then both compositions $\mathcal{L}_{1}\circ\mathcal{L}_{2}$ and $\mathcal{L}_{1}\circ\mathcal{L}_{2}$ are decomposable.
@@ 6. Optimal lower bounds for the Sindici–Piani semidefinite program
Furthermore, if $0<u<k$, one verifies that the four points $(a,b)$ again listed in Table 1 lie on the four distinct sides of this parallelogram. Therefore, the conic $\tilde{f}_{u}^{(d,k)}(x,y)=0$ defines an ellipse inscribed in the above parallelogram whenever $0<u<k$.

### v2

@@ 1. Introduction
The theory of positive linear maps plays a central role in operator algebras and quantum information theory [Stø63, Cho75, Wor76, Pau02, Stø13, Wat18]. Among the various refinements of positivity, the notion of $k$-positivity has emerged as a fundamental yet notoriously subtle concept. A linear map $\mathcal{L}:M_{d_{A}}(\mathbb{C})\to M_{d_{B}}(\mathbb{C})$ between matrix algebras is said to be $k$-positive if its ampliation map $\mathrm{id}_{k}\otimes\mathcal{L}:M_{k}(\mathbb{C})\otimes M_{d_{A}}(\mathbb{C})\to M_{k}(\mathbb{C})\otimes M_{d_{B}}(\mathbb{C})$ maps positive semidefinite operators to positive semidefinite operators. While complete positivity ($k=\min(d_{A},d_{B})$) admits several powerful characterizations, including the Choi–Kraus representation, the structure of $k$-positive maps for intermediate values of $k$ remains poorly understood. In fact, even deciding ($1$-)positivity is computationally hard [Gur03, GHP10], and several semidefinite programming relaxations for testing $k$-positivity have been proposed recently [CCF25]. Nevertheless, no general structural classification is known beyond very special cases.
@@
In the case of the $d\otimes d$ system, evidence for the existence of PPT states whose Schmidt number scales linearly in $d$ was first obtained in [SWZ11] using tools from asymptotic geometric analysis, although explicit constructions remained elusive. Early explicit examples exhibited PPT states whose Schmidt number grows logarithmically in $d$ [CYT17], while subsequent works improved this bound to linear scaling in $d$ [HLLMH18, PV19, Car20]. At present, the best known general upper bound is $\lfloor d/2\rfloor$, except for the dimensions $d=3$ [Cho82, Hor97] and $d=5$ [KG24], and explicit constructions achieving Schmidt number of order $d/2$ has been known only in limited settings [PV19, Car20]. The well-known dual formulation of this problem asks for the existence of $k$-positive indecomposable (i.e., not of the form $\mathcal{L}^{\prime}+\top\circ\mathcal{L}^{\prime\prime}$ for some completely positive maps $\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime\prime}$) linear maps on $M_{d}(\mathbb{C})$ as large as possible, a question originally raised in [Ter01]. This dual viewpoint has also driven much of the development of the theory of $k$-positive maps and its Choi matrices [CK09, YLT16, HLLMH18, RBO20, MOM25, HK25a, HK25b, vEKD25].
@@
In this work, we contribute to both of the above questions on $k$-positivity and Schmidt numbers by exploring new classes of quantum objects possessing symmetry with respect to the symplectic group. More precisely, we study linear maps that are covariant under conjugation by unitary symplectic matrices $S$, as well as bipartite quantum states that are invariant under $S\otimes S$ or $S\otimes\overline{S}$ actions, each parametrized by two real variables. The representation theory of the symplectic group gives rise to algebraic structures that are formally similar to those arising in symmetry classes associated with the orthogonal group [VW01, PY24]; however, the resulting entanglement properties turn out to be drastically different. As we show, this symmetry setting supports a remarkably rich entanglement landscape, including highly entangled PPT states and highly positive yet indecomposable linear maps.
@@
    in a $d\otimes d$ system, whenever it is well defined, is computed explicitly for all values of $(a,b)$. Furthermore, there exists a (triangular) set $S=S^{(d)}\subset\mathbb{R}^{2}$ with nonempty interior such that $\rho_{a,b}^{(d)}$ is ‘PPT with the Schmidt number $d/2$’ if and only if $(a,b)\in S$, while ${\rm SN}\big(\rho_{a,b}^{(d)}\big)\leq d/2$ whenever $\rho_{a,b}^{(d)}$ is PPT.
@@
  • When $p=q=-\frac{1}{d-2}$, the linear map $\mathcal{L}_{p,q}^{(d)}$ corresponds to the widely studied Breuer-Hall map [Bre06, Hal06] $\mathcal{L}^{\rm BH}$, which has proven effective in detecting PPT entanglement. It follows, however, that $\mathcal{L}^{\rm BH}$ fails to be $2$-positive; consequently, it cannot be used as an entanglement witness tailored solely to states with Schmidt number greater than $2$. In fact, $\mathcal{L}^{\rm BH}$ is shown to be atomic, that is, it is not of the form $\mathcal{L}^{\rm BH}=\mathcal{L}^{\prime}+\top\circ\mathcal{L}^{\prime\prime}$ for some $2$-positive linear maps $\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime\prime}$.
@@
    ${{\rm SN}}(\rho)-{\rm SN}(\rho^{\Gamma})=d/2-2$
    $=\max\{{{\rm SN}}(\rho_{a^{\prime},b^{\prime}}):\rho_{a^{\prime},b^{\prime}}\text{ is PPT}\}-\min\{\rm SN(\rho^{\prime}):\rho^{\prime}\text{ is entangled}\}.$
@@
In particular, we emphasize a new construction of Breuer–Hall–type maps, which are shown to be $k$-positive and indecomposable in strong sense and thus remedy a limitation of the original Breuer–Hall map.
Theorem 1.2 (Theorem 3.7, Remark 3.8).
For $d\geq 4$ an even integer and for $V$ a skew-symmetric unitary matrix, we consider the linear maps
@@
Then for $k=1,\ldots,d/2-1$,
  1. (1)
    $\mathcal{L}_{k}^{\rm BH}$ is an optimal $k$-positive indecomposable linear map: there exists no other $k$-positive witness that is able to detect strictly more PPT states with Schmidt number larger than $k$;
  2. (2)
    $\mathcal{L}^{\rm BH}_{k}$ cannot be decomposed as $\mathcal{L}^{\prime}+\top\circ\mathcal{L}^{\prime\prime}$ for any $(k+1)$-positive linear maps $\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime\prime}$.
@@
We have two additional applications using our framework. In Section 5, we show that the PPT squared conjecture [Chr12, CMHW19] holds within symplectic group symmetry in two perspectives: (1) any composition of two PPT maps is entanglement-breaking, and (2) any composition of a positive map and a PPT map is decomposable. This result is in fact a simple consequence of Theorem 1.1. Nevertheless, this is still remarkable in the sense that the verification applies to a broad class of quantum channels whose Choi matrices may exhibit high-dimensionality of entanglement, thereby providing new and nontrivial evidence in support of the conjecture.
@@ 2.3. Symplectic group symmetry
$V=\Omega_{d}:=\begin{pmatrix}0&I_{d/2}\\ -I_{d/2}&0\end{pmatrix};\quad\langle x,y\rangle_{\Omega}:=\sum_{i=1}^{d/2}(x_{i}y_{i+d/2}-x_{i+d/2}y_{i}),$
@@ 3.1. Positivity, complete positivity, and decomposability
Now suppose $\mathcal{L}_{p,q}$ is decomposable, and write $\mathcal{L}_{p,q}=\mathcal{L}^{(1)}+\mathcal{L}^{(2)}$ where $\mathcal{L}^{{}^{\prime}}$ is CP and $\mathcal{L}^{{}^{\prime\prime}}$ is co-CP. By applying the $(S,S)$-twirling operation $\mathcal{T}_{S,S}:=\mathcal{T}_{\iota_{Sp(d)},\iota_{Sp(d)}}$, we have $\mathcal{L}_{p,q}=\mathcal{T}_{S,S}\mathcal{L}^{{}^{\prime}}+\mathcal{T}_{S,S}{\mathcal{L}^{\prime\prime}}$ where $\mathcal{T}_{S,S}\mathcal{L}^{{}^{\prime}}$ and $\mathcal{T}_{S,S}\mathcal{L}^{{}^{\prime\prime}}$ are $(S,S)$-covariant CP and is co-CP maps, resp, by Proposition 2.5. Therefore, the trace-preserving property of $\mathcal{L}_{p,q}$, combined with Proposition 3.1, implies that
@@ 3.2. $k$-Positivity
It is straightforward to verify that the point $P=(-\frac{2}{d^{2}-3d},-\frac{d-2}{d^{2}-3d})$ satisfies the inequalities above, and hence $P\in\mathbb{P}_{d/2-1}\setminus\mathbb{D}$. On the other hand, the boundary line segment of $\mathbb{P}_{d/2}$ connecting the vertices $(-\frac{2}{d^{2}-d-2},-\frac{d}{d^{2}-d-2})$ and $(-\frac{2}{d^{2}-2},0)$ lies entirely within the half-plane $x+y\geq-\frac{2+d}{d^{2}-d-2}$. This implies $\mathbb{P}_{d/2}\subset\mathbb{D}$. ∎
@@
We recover the result of [Bre06] in the case $k=1$ in a concise manner by exploiting symplectic group symmetry. In fact, the same method used in the proof of Theorem 3.7 enables us to obtain further characterizations of linear maps of interest, as summarized below.
  1. (1)
    Without imposing the indecomposability condition, the optimal $k$-positive linear maps within ${\rm Cov}(S,S)$ corresponds precisely to, up to positive multiplicative constant,
    $\textstyle\{\mathcal{L}_{p,q}:(1-d)p+q=1,\;\;\frac{-1}{d-2}\leq p\leq 0\}$ for $k=1$,
    $\textstyle\{\mathcal{L}_{p,q}:(1-kd)p+q=1,\;\;\frac{-1}{kd-k-1}\leq p\leq\frac{-1}{kd-1}\}$ for $1<k<d$ and $k$ even,
    $\{\mathcal{L}_{k}^{\rm BH}\}\cup\{\mathcal{L}_{p,q}:(p,q)\in\partial\,\mathbb{P}_{k}\setminus\mathbb{P}_{d},\;\;p+(1-d)q\neq 1\}$ for $1<k<d$ and $k$ odd,
    where $\partial\,\mathbb{P}_{k}$ denotes the boundary of $\mathbb{P}_{k}$. For example, the $k$-reduction map
@@
    is optimal in the sense that there exists no nonzero CP map $\Phi$ such that $\mathcal{L}^{\prime}=\mathcal{L}_{k}^{\rm red}-\Phi$ remains $k$-positive, as also established in [SBL01].
  2. (2)
[…]
    $p+q<-\frac{k+2}{(k+1)d-k-2},\quad p+(1-d)q\leq 1,\quad(1-kd)p+q\leq 1,$
    or equivalently $(p,q)\in\mathbb{P}_{k}\setminus{\rm conv}(\mathbb{P}_{k+1}\cup\mathbb{P}_{k+1}^{\top})$, where $\mathbb{P}_{k+1}^{\top}=\{(p,q):(q,p)\in\mathbb{P}_{k+1}\}$ denotes the reflection of $\mathbb{P}_{k+1}$ across the line $y=x$. Note that the $k$-Breuer-Hall map $\mathcal{L}_{k}^{\rm BH}$ satisfies this property, and the set $\mathbb{P}_{k}\setminus{\rm conv}(\mathbb{P}_{k+1}\cup\mathbb{P}_{k+1}^{\top})$ has nonempty interior. When $k=1$, this property is known as atomicity. In particular, the Breuer–Hall map is atomic, recovering the result of [CK08].
@@ 5. The PPT squared conjecture under symplectic group symmetry
    If $\mathcal{L}_{1}$ is a positive map that is either $(S,S)$- or $(S,\overline{S})$-covariant, and $\mathcal{L}_{2}$ is a PPT map that is either $(S,S)$- or $(S,\overline{S})$-covariant, then both compositions $\mathcal{L}_{1}\circ\mathcal{L}_{2}$ and $\mathcal{L}_{1}\circ\mathcal{L}_{2}$ are decomposable.
@@ 6. Optimal lower bounds for the Sindici–Piani semidefinite program
Furthermore, if $0<u<k$, one verifies that the four points $(a,b)$ listed in Table 1 lie on the four distinct sides of this parallelogram. Therefore, the conic $\tilde{f}_{u}^{(d,k)}(x,y)=0$ defines an ellipse inscribed in the above parallelogram whenever $0<u<k$.